# Rigidity for von Neumann algebras of graph product groups. I. Structure of automorphisms

### v2

@@ Abstract
In this paper we study various rigidity aspects of the von Neumann algebra $L(\Gamma)$ where $\Gamma$ is a graph product group whose underlying graph is a certain cycle of cliques and the vertex groups are wreath-like product property (T) groups. Using an approach that combines methods from Popa’s deformation/rigidity theory with new techniques pertaining to graph product algebras, we describe all symmetries of these von Neumann algebras and reduced $*{}^{*}$-algebras by establishing formulas in the spirit of Genevois and Martin’s results on automorphisms of graph product groups.
@@ 1.1 Statements of the main results
[…]
@@
We conclude our introduction by describing in Corollary D all ***-isomorphisms of the reduced $*{}^{*}$-algebras of graph product groups that we considered in Theorem B. This result can be seen as a $*{}^{*}$-algebraic version of Genevois and Martin’s result [GM19, Corollary C].
@@
In fact, this result is a consequence of Theorem B since the graph product groups that we consider have trivial amenable radical (see Lemma 4.3), and consequently, their reduced $*{}^{*}$-algebras have unique trace [BKKO14].
Acknowledgements. We would like to thank Stefaan Vaes for helpful comments. We would also like to thank the referee for all the comments and suggestions that greatly improved the exposition of the paper.
@@ 2.3 Popa’s intertwining-by-bimodules techniques
    There exist projections $p_{0}\in\mathcal{P},q_{0}\in\mathcal{Q}$, a ***-homomorphism $\theta:p_{0}\mathcal{P}p_{0}\rightarrow q_{0}\mathcal{Q}q_{0}$ and a non-zero partial isometry $v\in q_{0}\mathcal{M}p_{0}$ such that $\theta(x)v=vx$, for all $x\in p_{0}\mathcal{P}p_{0}$.
@@ 2.5 A result on normalizers in tensor product factors
[…]
@@
[…]
@@
[…]
@@ 3.1 Unique prime factorization for von Neumann algebras of wreath-like product groups
[…]
@@ 4 Graph product groups associated with cycles of cliques graphs
[…]
@@
[…]
@@ 7 Symmetries of graph product group von Neumann algebras
[…]
@@
There exists a unique normal-$\ast$-isomorphism denoted by $(\Phi,\sigma):\mathcal{L}(\Gamma_{\mathscr{G}})\rightarrow\mathcal{L}(\Lambda_{\mathscr{H}})$ which extends the maps $\bigcup_{v\in\mathscr{V}}\mathcal{L}(\Gamma_{v})\ni x\rightarrow\Phi^{\sigma}_{v}(x)\in\mathcal{L}(\Lambda_{\mathscr{H}})$.
@@
Let $\Gamma=\mathscr{G}\{\Gamma_{v}\}$ and $\Lambda=\mathscr{H}\{\Lambda_{w}\}$ be graph product groups and assume that $\Gamma_{v}$ and $\Lambda_{w}$ are icc property (T) groups for all $v\in\mathscr{V}$ and $w\in\mathscr{W}$. Let $\theta:\mathcal{L}(\Gamma)\rightarrow\mathcal{L}(\Lambda)$ be any normal-$\ast$-isomorphism. Then there is a bijection $\sigma:{\rm cliq}(\mathscr{G})\rightarrow{\rm cliq}(\mathscr{H})$ such that for every $\mathscr{C}\in{\rm cliq}(\mathscr{G})$ there is a unitary $u_{\mathscr{C}}\in\mathcal{L}(\Lambda)$ such that $\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))=u_{\mathscr{C}}\mathcal{L}(\Lambda_{\sigma(\mathscr{C})})u_{\mathscr{C}}^{*}$.
@@
Reversing the roles of $\Gamma$ and $\Lambda$, in a similar manner for every $\mathscr{D}\in{\rm cliq}(\mathscr{H})$ one can find $\tau(\mathscr{D})\in{\rm cliq}(\mathscr{G})$ and a unitary $w_{\mathscr{D}}\in\mathcal{L}(\Gamma)$ satisfying $*.\mathcal{L}(\Lambda_{\mathscr{D}})\subseteq w_{\mathscr{C}}\theta(\mathcal{L}(\Gamma_{\tau(\mathscr{D})})w^{*}_{\mathscr{D}}.$ Altogether these show that $*u_{\mathscr{C}}\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))u^{*}_{\mathscr{C}}\subseteq\mathcal{L}(\Lambda_{\sigma(\mathscr{C})})\subseteq w_{\sigma(\mathscr{C})}\theta(\mathcal{L}(\Gamma_{\tau(\sigma(\mathscr{C}))})w^{*}_{\sigma(\mathscr{C})}$. In particular, Theorem 2.7 implies that $\mathscr{C}=\tau(\sigma(\mathscr{C}))$ and $u_{\mathscr{C}}^{*}w_{\sigma(\mathscr{C})}\in\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))$. This combined with the prior containment imply that $u_{\mathscr{C}}\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))u^{*}_{\mathscr{C}}=\mathcal{L}(\Lambda_{\sigma(\mathscr{C})})$. As $\mathscr{C}=\tau(\sigma(\mathscr{C}))$ for any clique $\mathscr{C}$ of $\mathscr{G}$, it follows in particular that $\sigma$ is a bijection. ∎
@@
Let $10<t<$ 1 be a scalar and $\Theta:\mathcal{L}(\Gamma)^{t}\rightarrow\mathcal{L}(\Lambda)$ be any normal-$\ast$-isomorphism.
@@
[…]
@@
[…]
@@ 7.3 Proofs of the main results
Let $t>0$ and let $\Theta:\mathcal{L}(\Gamma)^{t}\rightarrow\mathcal{L}(\Lambda)$ be any normal-$\ast$-isomorphism. Then $t=1$ and one can find an isometry $\sigma:\mathscr{G}\rightarrow\mathscr{H}$, normal-$\ast$-isomorphisms $\theta_{i-1,i}:\mathcal{L}(\Gamma_{\mathscr{C}_{i-1,i}})\rightarrow\mathcal{L}(\Gamma_{\mathscr{C}_{\sigma(\mathscr{C}_{i-1,i})}})$, $\xi_{i}:\mathcal{L}(\Gamma_{\mathscr{C}^{\rm int}_{i}})\rightarrow\mathcal{L}(\Gamma_{\sigma(\mathscr{C}^{\rm int}_{i})})$ for all $i\in\overline{1,n}$, and a unitary $u\in\mathcal{L}(\Lambda)$ such that $\Theta={\rm ad}(u)\circ\phi_{\theta,\xi}$.
@@
Using the $*{}^{*}$-superrigid property (T) wreath-like product groups recently discovered in [CIOS21] as vertex groups in the previous result one obtains an even more precise description of the $\ast$-isomorphisms between these von Neumann algebras; hence, we provide a proof for Theorem B.
@@
Then for any $t>0$ and normal-$\ast$-isomorphism $\Theta:\mathcal{L}(\Gamma)^{t}\rightarrow\mathcal{L}(\Lambda)$ we have $t=1$ and one can find a character $\eta\in{\rm Char}(\Gamma)$, a group isomorphism $\delta\in{\rm Isom}(\Gamma,\Lambda)$, an automorphism of $\mathcal{L}(\Lambda)$ of the form $\phi_{a,b}$ (see the notation after equation (7.1)) and a unitary $u\in\mathcal{L}(\Lambda)$ such that $\Theta={\rm ad}(u^{*})\circ\phi_{a,b}\circ\Psi_{\eta,\delta}$.
@@
Then for any normal-$\ast$-isomorphism $\Theta:C_{r}^{*}(\Gamma)\rightarrow C_{r}^{*}(\Lambda)$, one can find a character $\eta\in{\rm Char}(\Gamma)$, a group isomorphism $\delta\in{\rm Isom}(\Gamma,\Lambda)$, an automorphism of $\mathcal{L}(\Lambda)$ of the form $\phi_{a,b}$ and a unitary $u\in\mathcal{L}(\Lambda)$ such that $\Theta={\rm ad}(u^{*})\circ\phi_{a,b}\circ\Psi_{\eta,\delta}$.
@@
From Lemma 4.3 we get that $\Gamma$ has trivial amenable radical, and hence, by [BKKO14, Theorem 1.3] it follows that $C_{r}^{*}(\Gamma)$ has unique trace. This implies that any ***-isomorphism between $C_{r}^{*}(\Gamma)$ and $C_{r}^{*}(\Lambda)$ lifts to a ***-isomorphism of the associated von Neumann algebras. Now the result follows from Theorem 7.10. ∎

### v3

@@ Abstract
In this paper we study various rigidity aspects of the von Neumann algebra $L(\Gamma)$ where $\Gamma$ is a graph product group whose underlying graph is a certain cycle of cliques and the vertex groups are wreath-like product property (T) groups. Using an approach that combines methods from Popa’s deformation/rigidity theory with new techniques pertaining to graph product algebras, we describe all symmetries of these von Neumann algebras and reduced C-algebras by establishing formulas in the spirit of Genevois and Martin’s results on automorphisms of graph product groups.
@@ 1.1 Statements of the main results
This theorem applies to fairly large classes of property (T) vertex groups, including: all fibered Rips constructions considered in [CDK19, CDHK20] and all wreath-like product groups $\mathcal{W}\mathcal{R}(A,B\curvearrowright I)$ where $A$ is either abelian or icc, $B$ is an icc subgroup of a hyperbolic group, and the action $B\curvearrowright I$ has amenable stabilizers, [CIOS21]. The result also implies that the fundamental group [MvN36] of these graph product group II1 factors is always trivial; this means that if $\Gamma$ is a graph product group as in Theorem A, then $\{t>0|\;\mathcal{L}(\Gamma)^{t}\cong\mathcal{L}(\Gamma)\}=1$. Recall that Popa used his deformation/rigidity theory for obtaining the first examples of II1 factors with trivial fundamental group [Po01], hence answering a longstanding open problem of Kadison, see [Ge03]. Subsequently, a large number of striking results on computations of fundamental groups of II1 factors were obtained, see the introduction of [CDHK20]. To our knowledge, Theorem A provides the first instance of computing the fundamental group for non-trivial graph product von Neumann algebras which is not a tensor product.
@@
We conclude our introduction by describing in Corollary D all *-isomorphisms of the reduced C-algebras of graph product groups that we considered in Theorem B. This result can be seen as a C-algebraic version of Genevois and Martin’s result [GM19, Corollary C].
@@
In fact, this result is a consequence of Theorem B since the graph product groups that we consider have trivial amenable radical (see Lemma 4.3), and consequently, their reduced C-algebras have unique trace [BKKO14].
Acknowledgements. We would like to thank Stefaan Vaes for helpful comments. We would also like to thank the referee for all the comments and suggestions that greatly improved the exposition of the paper. I.C. was partially supported by NSF FRG Grant #1854194 and NSF Grant #DMS-2154637; D.D. was supported by the postdoctoral fellowship fundamental research 12T5221N of the Research Foundation Flanders.
@@ 2.3 Popa’s intertwining-by-bimodules techniques
    There exist projections $p_{0}\in\mathcal{P},q_{0}\in\mathcal{Q}$, a *-homomorphism $\theta:p_{0}\mathcal{P}p_{0}\rightarrow q_{0}\mathcal{Q}q_{0}$ and a non-zero partial isometry $v\in q_{0}\mathcal{M}p_{0}$ such that $\theta(x)v=vx$, for all $x\in p_{0}\mathcal{P}p_{0}$.
@@ 2.5 A result on normalizers in tensor product factors
Our next proposition describes the normalizer of a II1 factor $\mathcal{N}$ inside the tensor product of $\mathcal{N}$ with another II1 factor. More precisely, we have:
@@
Let $\mathcal{N}$ and $\mathcal{P}$ be II1 factors and denote by $\mathcal{M}=\mathcal{N}\bar{\otimes}\mathcal{P}$. If $u\in\mathscr{U}(\mathcal{M})$ satisfies $u\mathcal{N}u^{*}=\mathcal{N}$, then one can find $a\in\mathscr{U}(\mathcal{N})$ and $b\in\mathscr{U}(\mathcal{P})$ such that $u=a\otimes b$.
@@
Let $u_{i}\in\mathcal{N}$ be the partial isometry in the polar decomposition of $\eta_{i}$. Thus $\theta(x)u_{i}=u_{i}x$ for all $x\in\mathcal{N}$ and all $i$. In particular we get $u^{*}_{i}u_{i}\in\mathcal{N}^{\prime}\cap\mathcal{N}=\mathbb{C}1$ and hence $u_{i}\in\mathscr{U}(\mathcal{N})$ for all $i$. The prior relations also imply that $u_{i}^{*}xu_{i}=\theta(x)=u_{j}^{*}xu_{j}$ for all $i,j\in I$. In particular, we have that $u_{i}u_{j}^{*}\in\mathcal{N}^{\prime}\cap\mathcal{N}=\mathbb{C}1$ and thus one can find scalars $c_{i,j}\in\mathbb{T}$ so that $u_{i}=c_{i,j}u_{j}$ for all $i,j\in I$. Relation (2.2) also implies that $|\eta_{i}|\in\mathcal{N}^{\prime}\cap L^{2}(\mathcal{N})$ and since $\mathcal{N}$ is a II1 factor we get that $|\eta_{i}|\in\mathbb{C}1$. In conclusion, $\eta_{i}\in\mathbb{C}\mathscr{U}(\mathcal{N})$ for all $i$ and one can find $d_{i,j}\in\mathbb{C}$ such that $\eta_{i}=d_{i,j}\eta_{j}$ for all $i,j\in I$. Fix $j\in I$ with $\eta_{j}\neq 0$. Using the above relations we have that $u=\sum_{i}\eta_{i}\otimes\xi_{i}=\sum_{i}d_{i,j}\eta_{j}\otimes\xi_{i}=\eta_{j}\otimes(\sum_{i}d_{i,j}\xi_{i})=\eta_{j}\otimes b$, where we denoted by $b=\sum_{i}d_{i,j}\xi_{i}\in L^{2}(\mathcal{P})$. Since $\eta_{j}\in\mathbb{C}\mathscr{U}(\mathcal{N})$ we get the desired conclusion. ∎
@@ 3.1 Unique prime factorization for von Neumann algebras of wreath-like product groups
If $\mathcal{M}:=\mathcal{L}(\Gamma_{1})\bar{\otimes}\dots\bar{\otimes}\mathcal{L}(\Gamma_{n})=\mathcal{P}_{1}\bar{\otimes}\mathcal{P}_{2}$ is a tensor product decomposition into II1 factors, then there exist a unitary $u\in\mathcal{M}$, a decomposition $\mathcal{M}=\mathcal{P}_{1}^{t}\bar{\otimes}\mathcal{P}_{2}^{1/t}$ and a partition $T_{1}\sqcup T_{2}=\overline{1,n}$ such that (×$\mathcal{L}(\times_{k\in S_{j}}\Gamma_{k})=u\mathcal{P}_{j}^{t_{j}}u^{*}$, for any $j\in\{1,2\}.$
@@ 4 Graph product groups associated with cycles of cliques graphs
A basic example of a graph in the class CC1 is a simple cycle of triangles called $\mathcal{T}_{n}$, where $n$ is the number of cliques; see picture below for $n=16$.
@@
One can easily check that the graphs in CC1 are atomic and therefore the conclusion follows immediately from [GM19, Corollary C].∎
@@ 7 Symmetries of graph product group von Neumann algebras
The main result of this section is a strong rigidity result describing all $\ast$-isomorphisms between factors associated with a fairly large family of graph product groups arising from finite graphs in the class CC1 (Theorem 7.10). As a by-product we obtain concrete descriptions of all symmetries of these factors including such examples with trivial fundamental groups (Corollaries 7.11 and 7.12). However, to be able to state and prove these results we first need to introduce some new terminology and establish a few preliminary results.
@@
There exists a unique $\ast$-isomorphism denoted by $(\Phi,\sigma):\mathcal{L}(\Gamma_{\mathscr{G}})\rightarrow\mathcal{L}(\Lambda_{\mathscr{H}})$ which extends the maps $\bigcup_{v\in\mathscr{V}}\mathcal{L}(\Gamma_{v})\ni x\rightarrow\Phi^{\sigma}_{v}(x)\in\mathcal{L}(\Lambda_{\mathscr{H}})$.
@@
Let $\Gamma=\mathscr{G}\{\Gamma_{v}\}$ and $\Lambda=\mathscr{H}\{\Lambda_{w}\}$ be graph product groups and assume that $\Gamma_{v}$ and $\Lambda_{w}$ are icc property (T) groups for all $v\in\mathscr{V}$ and $w\in\mathscr{W}$. Let $\theta:\mathcal{L}(\Gamma)\rightarrow\mathcal{L}(\Lambda)$ be any $\ast$-isomorphism. Then there is a bijection $\sigma:{\rm cliq}(\mathscr{G})\rightarrow{\rm cliq}(\mathscr{H})$ such that for every $\mathscr{C}\in{\rm cliq}(\mathscr{G})$ there is a unitary $u_{\mathscr{C}}\in\mathcal{L}(\Lambda)$ such that $\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))=u_{\mathscr{C}}\mathcal{L}(\Lambda_{\sigma(\mathscr{C})})u_{\mathscr{C}}^{*}$.
@@
Reversing the roles of $\Gamma$ and $\Lambda$, in a similar manner for every $\mathscr{D}\in{\rm cliq}(\mathscr{H})$ one can find $\tau(\mathscr{D})\in{\rm cliq}(\mathscr{G})$ and a unitary $w_{\mathscr{D}}\in\mathcal{L}(\Gamma)$ satisfying $\mathcal{L}(\Lambda_{\mathscr{D}})\subseteq w_{\mathscr{C}}\theta(\mathcal{L}(\Gamma_{\tau(\mathscr{D})})w^{*}_{\mathscr{D}}.$ Altogether these show that $u_{\mathscr{C}}\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))u^{*}_{\mathscr{C}}\subseteq\mathcal{L}(\Lambda_{\sigma(\mathscr{C})})\subseteq w_{\sigma(\mathscr{C})}\theta(\mathcal{L}(\Gamma_{\tau(\sigma(\mathscr{C}))})w^{*}_{\sigma(\mathscr{C})}$. In particular, Theorem 2.7 implies that $\mathscr{C}=\tau(\sigma(\mathscr{C}))$ and $u_{\mathscr{C}}^{*}w_{\sigma(\mathscr{C})}\in\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))$. This combined with the prior containment imply that $u_{\mathscr{C}}\theta(\mathcal{L}(\Gamma_{\mathscr{C}}))u^{*}_{\mathscr{C}}=\mathcal{L}(\Lambda_{\sigma(\mathscr{C})})$. As $\mathscr{C}=\tau(\sigma(\mathscr{C}))$ for any clique $\mathscr{C}$ of $\mathscr{G}$, it follows in particular that $\sigma$ is a bijection. ∎
@@
Let $10<t<$ 1 be a scalar and $\Theta:\mathcal{L}(\Gamma)^{t}\rightarrow\mathcal{L}(\Lambda)$ be any $\ast$-isomorphism.
@@
Let $\mathscr{D}$ be a clique in $\mathscr{G}$. Since $\mathcal{L}(\Gamma_{\mathscr{D}})$ is a II1 factor, there is a projection $p\in\mathcal{L}(\Gamma_{\mathscr{D}})$ of trace $\tau(p)=t$ with $\mathcal{L}(\Gamma)^{t}=p\mathcal{L}(\Gamma)p$. As $\mathcal{L}(\Gamma_{\mathscr{D}})$ has property (T) then so is $p\mathcal{L}(\Gamma_{\mathscr{D}})p$. Since $p\mathcal{L}(\Gamma_{\mathscr{D}})p\subset\Theta^{-1}(\mathcal{L}(\Lambda)):=\mathcal{N}$ then by Theorem 6.1 one can find a clique $\mathscr{F}\in{\rm cliq}(\mathscr{H})$ such that $p\mathcal{L}(\Gamma_{\mathscr{D}})p\prec_{\mathcal{N}}\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}}))$. Now, observe that since the inclusion $p\mathcal{L}(\Gamma_{\mathcal{D}})p\subset\mathcal{N}$ is irreducible, we can proceed as in the proof of Theorem 7.5 to deduce that $p\mathcal{L}(\Gamma_{\mathscr{D}})p\nprec_{\mathcal{N}}\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}\setminus\{c\}}))$ for every $c\in\mathscr{F}$. Thus, using the irreducibility condition and the moreover part of Theorem 6.1, there is $u\in\mathscr{U}(\Theta^{-1}(\mathcal{L}(\Lambda)))$ satisfying
@@
Also observe that $(up\mathcal{L}(\Gamma_{\mathscr{D}})pu^{*})^{\prime}\cap\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}}))\subseteq up(\mathcal{L}(\Gamma_{\mathscr{D}}))^{\prime}{\cap\mathcal{L}(\Gamma_{\mathscr{G}})pu^{*}}=\mathbb{C}p$. Hence, (7.2) is an irreducible inclusion of II1 factors.
Next, since $\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}}))$ has property (T) and $\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}}))\subset p\mathcal{L}(\Gamma)p\subset\mathcal{L}(\Gamma):=\mathcal{M}$ then by Theorem 6.1 one can find $\mathscr{D}^{\prime}\in{\rm cliq}(\mathscr{G})$ such that $\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}}))\prec_{\mathcal{M}}\mathcal{L}(\Gamma_{\mathscr{D}^{\prime}})$. Combining this with (7.2) we further get $p\mathcal{L}(\Gamma_{\mathscr{D}})p\prec_{\mathcal{M}}\mathcal{L}(\Gamma_{\mathscr{D}^{\prime}})$, which further implies by Lemma 2.3 that $\mathscr{D}\subseteq\mathscr{D}^{\prime}$ and since these are cliques we conclude that $\mathscr{D}=\mathscr{D}^{\prime}$. In conclusion, the prior intertwining relation amounts to $\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}}))\prec_{\mathcal{M}}\mathcal{L}(\Gamma_{\mathscr{D}})$. Since $\mathcal{L}(\Gamma_{\mathscr{D}})$ is a II1 factor we further obtain $\Theta^{-1}(\mathcal{L}(\Lambda_{\mathscr{F}}))\prec_{\mathcal{M}}up\mathcal{L}(\Gamma_{\mathscr{D}})pu^{*}$. Since $u\in p\mathcal{M}p$ is a unitary this further implies that
@@ 7.3 Proofs of the main results
Let $t>0$ and let $\Theta:\mathcal{L}(\Gamma)^{t}\rightarrow\mathcal{L}(\Lambda)$ be any $\ast$-isomorphism. Then $t=1$ and one can find an isometry $\sigma:\mathscr{G}\rightarrow\mathscr{H}$, $\ast$-isomorphisms $\theta_{i-1,i}:\mathcal{L}(\Gamma_{\mathscr{C}_{i-1,i}})\rightarrow\mathcal{L}(\Gamma_{\mathscr{C}_{\sigma(\mathscr{C}_{i-1,i})}})$, $\xi_{i}:\mathcal{L}(\Gamma_{\mathscr{C}^{\rm int}_{i}})\rightarrow\mathcal{L}(\Gamma_{\sigma(\mathscr{C}^{\rm int}_{i})})$ for all $i\in\overline{1,n}$, and a unitary $u\in\mathcal{L}(\Lambda)$ such that $\Theta={\rm ad}(u)\circ\phi_{\theta,\xi}$.
@@
Using the W-superrigid property (T) wreath-like product groups recently discovered in [CIOS21] as vertex groups in the previous result one obtains an even more precise description of the $\ast$-isomorphisms between these von Neumann algebras; hence, we provide a proof for Theorem B.
@@
Then for any $t>0$ and $\ast$-isomorphism $\Theta:\mathcal{L}(\Gamma)^{t}\rightarrow\mathcal{L}(\Lambda)$ we have $t=1$ and one can find a character $\eta\in{\rm Char}(\Gamma)$, a group isomorphism $\delta\in{\rm Isom}(\Gamma,\Lambda)$, an automorphism of $\mathcal{L}(\Lambda)$ of the form $\phi_{a,b}$ (see the notation after equation (7.1)) and a unitary $u\in\mathcal{L}(\Lambda)$ such that $\Theta={\rm ad}(u^{*})\circ\phi_{a,b}\circ\Psi_{\eta,\delta}$.
@@
Then for any $\ast$-isomorphism $\Theta:C_{r}^{*}(\Gamma)\rightarrow C_{r}^{*}(\Lambda)$, one can find a character $\eta\in{\rm Char}(\Gamma)$, a group isomorphism $\delta\in{\rm Isom}(\Gamma,\Lambda)$, an automorphism of $\mathcal{L}(\Lambda)$ of the form $\phi_{a,b}$ and a unitary $u\in\mathcal{L}(\Lambda)$ such that $\Theta={\rm ad}(u^{*})\circ\phi_{a,b}\circ\Psi_{\eta,\delta}$.
@@
From Lemma 4.3 we get that $\Gamma$ has trivial amenable radical, and hence, by [BKKO14, Theorem 1.3] it follows that $C_{r}^{*}(\Gamma)$ has unique trace. This implies that any *-isomorphism between $C_{r}^{*}(\Gamma)$ and $C_{r}^{*}(\Lambda)$ lifts to a *-isomorphism of the associated von Neumann algebras. Now the result follows from Theorem 7.10. ∎